Is the function f, defined by f(x)= x^2-5x+4/(x^2-16) , x≠4

f(4)=0, continuous at x=4. Give reasons for your answer.


1
Expert's answer
2021-05-28T09:56:30-0400

The function f is continuous at some point c if:

"\displaystyle {\lim_{x\to c} }=f(c)"

In our case:

"\displaystyle {\lim_{x\to 4} }\ne f(4)"

So, the function f(x) is not continuous at x=4